Find the distance between (-3,2) and (1,4). Please show work.
Answer by Alan3354(69443)

Find the distance between (-3,2) and (1,4). Please show work.
---------------------------
Use Pythagoras. The distance squared is the (diff in x)^2 plus the (diff in y)^2
s^2 = (x2-x1)^2 + (y2-y1)^2
s^2 = (1 -(-3))^2 + (4-2)^2
= (4)^2 + (2)^2
= 16 + 4
= 20
s = sqrt(20) = 2sqrt(5)